# OODTE: A Differential Testing Engine for the ONNX Optimizer

### v3

@@ 3.5. Fault Localization
Once a method is selected, the Metrics Comparison Module calculates the metrics based on model type and generates a report with all results, in JSON format. If a difference is detected across the original and the optimized model, then OODTE retrofits the reported data to the Optimizer Module. Then, the Optimizer Module generates model optimizations per-pass, reporting how each pass behaves. OODTE applies a fault localization methodology for three cases: when the optimizer crashes, when it generates an optimized model in an invalid form, and when the optimization succeeds but produces results inconsistent with the original model. In addition, OODTE registers any warnings that can indicate the presence of a potential performance fault (e.g., redundant initializers in the model). Upon detecting an issue or a problematic behavior, OODTE attempts to identify the responsible pass by generating per-pass optimizations, executing inference on the resulting models, and comparing their outputs with those of the original model. This process is applied iteratively for each of the passes supported by the optimizer, until the pass responsible for the detected issue is identified.
@@ 4. Related Work
Table 1. Comparison of features between OODTE and related work. Each technique is compared across the following parameters: testing of ONNX (ONNX?), ability for fault localization (Flt. Loc.), bug types, graph optimization passes (Gr. Opt. Passes), detection of accuracy changes (Acc.), crashes (Cr.), generation of malformed models due to miscompilation (Malf.), and issues affecting model execution times(T.) bugs, as well as the compilers explored on each contribution.
@@
OODTE
Per-pass fault
localization,
differential
testing
Logic, API versioning,
graph input/parameters/
transformations
Fuse/
eliminate/
rewrite
ONNX
Optimizer (onnxoptimizer, )
NNSmith (NNSmith, )
Grammar-
based fuzzing
†,‡ $\times$
Conversions, semantics,
gen.transformations,
graph rewrites
Conv2D
inputs
rewrite
$\times$ $\times$
TVM (tvm, ),
PyTorch JIT (pytorch, ),
TensorRT (tensorrt, ),
PT2 (pytorch2, ), ONNX
Runtime (onnxruntime, ),
TF XLA (tfxla, ),
TFLite (tensorflow, )
MLIRSmith (MLIRSmith, )
Grammar-
based fuzzing
$\times$ $\times$
Conversions (dialects),
generic transformations,
logic (dialects)
$\times$ $\times$ $\times$ $\times$
MLIR (MLIR, ),
LLVM (LLVM, )
MLIRod (SuoMLIR, )
Operator
dependency
analysis/
fuzzing
$\times$ $\times$
Conversions (dialects),
gen. transformations
$\times$ $\times$ $\times$ $\times$ MLIR (MLIR, )
SYNTH-
FUZZ (synthfuzz, )
Grammar-
based fuzzing
@@
$\times$ $\times$ Generic, input-based $\times$ $\times$ $\times$ $\times$ MLIR (MLIR, )
HirGen (HirGen, )
Graph IR
fuzzing
$\times$ $\times$
Numeric computations,
exceptions, types, memory
$\times$ $\times$ $\times$ $\times$ TVM (tvm, )
TVMFuzz (TVMFuzz, )
Low-level
IR fuzzing
$\times$ $\times$
Numeric computations,
graph rewrites
Constant
folding
$\times$ $\times$ TVM (tvm, )
Tzer (TZER, )
Coverage-
guided fuzzing
(Low-level IR)
$\times$ $\times$
API inconsistency/misuse,
memory, types
$\times$ $\times$ TVM (tvm, )
GenCoG (GenCoG, )
Constraint
solving,
High-level
IR Fuzzing
$\times$ $\times$
Memory, tensor shape
/data, operator attributes,
logic
Operator
fusion
$\times$ $\times$ TVM (tvm, )
DeepDiffer (DeepDiffer, )
Low-level IR/
coverage-
guided fuzzing,
differential
testing
$\times$ $\times$
Type, memory, API misuse,
numeric computations,
exception handling
$\times$ $\times$ $\times$ TVM (tvm, )
WhiteFox (WhiteFox, )
LM analysis,
fuzzing
$\times$ $\times$
Memory, API misuse,
concurrency, logic,
security vulnerabilities
$\times$ $\times$
TF XLA (tfxla, ),
PyTorch
Inductor (pytorch, ),
LLVM (LLVM, )
TorchProbe (Su_2023, )
Feature-
based fuzzing
$\times$ $\times$
Memory, API misuse,
logic
$\times$ $\times$ $\times$
PyTorch (pytorch, ),
Triton (TilletTriton, )
MT-
DLComp (Xiao_2022, )
Metamorphic
testing
$\times$ $\times$
Memory, dead code,
layout optimization,
numeric accuracy
Conv2D
layout
change
$\times$ $\times$
TVM (tvm, ),
Glow (Glow, ),
NNFusion (NNFusion, ),
TF XLA (tfxla, )
FetaFix (FetaFix, )
Fault localiza-
tion,
differential
testing
$\times$
Preprocessing, input
dimensions, parameters,
hyperparameters, tensor
shape, graph structure
$\times$ $\times$ $\times$ $\times$
Converters across
PyTorch (pytorch, ),
Keras (chollet2015keras, ),
TF (tensorflow, ), TFLite (tensorflow, )
In terms of the Technique column shown in Table 1, most existing works primarily rely on compiler fuzzing, often enhanced with techniques such as differential testing (e.g., DeepDiffer (DeepDiffer, )) or LLM analysis (e.g., WhiteFox (WhiteFox, )). Consequently, these systems focus on triggering compiler bugs rather than pinpointing their origins and root causes, typically relying on generic compiler error messages to perform manual fault localization. Among these works, only FetaFix (FetaFix, ) provides a mechanism for fault localization, but it targets accuracy-related faults in model conversion between DL frameworks and does not support the model optimization process. In contrast, OODTE systematically localizes faults by applying an iterative fault localization methodology for each pass in the ONNX Optimizer.
In terms of bug types detected in Table 1, existing tools primarily focus on detecting crash bugs. For label accuracy-related bugs, they typically rely on random input generation (e.g., NNSmith (NNSmith, )) to expose inconsistencies in model inference. In contrast, for label inconsistency bugs that OODTE detects, we assess the prevalence and impact of these bugs on popular datasets(e.g., ILSVRC (ILSVRC17, ), IMDB (imdbsentiment, )) and model architectures. Additionally, only a small number of tools consider miscompilations (for instance, models that contain problematic input nodes or misconfigured parameters), with just two prior works capable of detecting faults of this kind (TVMFuzz, ; WhiteFox, ). The detection of bugs affecting execution times of the compiled models has received limited attention in the literature, with only two TVM-focused systems (TZER, ; DeepDiffer, ) using empirical measurements to evaluate performance-related faults. OODTE currently identifies issues that could potentially lead to degraded model execution times by flagging redundant operator warnings (e.g., unused initializer entries), while the integration of precise execution time measurements is left as subject for future work.
Regarding support for ONNX in Table 1, OODTE is the only tool capable of localizing faults in the ONNX Optimizer. Other methodologies either treat ONNX merely as an intermediate representation or focus on bug detection in ONNX Runtime (WhiteFox, ), a cross-platform inference engine, unlike model optimization targeted by the ONNX Optimizer.
Existing work on bug detection within graph optimization passes (shown as Gr. Opt. Passes column in Table 1) is limited, with prior approaches focusing on common operations such as constant folding, operator fusion, and convolution optimizations. In contrast, OODTE addresses the full spectrum of features in the ONNX Optimizer, covering a set of 47 passes that perform a wide variety of graph optimizations such as Fuse, Eliminate, and Rewrite operations in addition to the more common operations.
@@ 5.1. The Experiment Process
We evaluated all ONNX Optimizer passes (Section 2.3). For each model, we first applied the default setting, which runs all fusion and elimination passes, and compared the outputs of the original and optimized models as described in Section 3.4. When discrepancies arose, we re-optimized using individual passes, producing 47 single-pass models to identify the pass causing the issue. OODTE also supports new or custom passes, and our methodology generalizes to other optimization frameworks (e.g., Apache TVM) that allow both bundled and individual pass execution.
@@ 5.2. DNN Models
For our experiment set, we sourced $130$ model instances from the ONNX Model Hub (onnxmodelhub, ). In particular, we selected $93$ models utilized for classification tasks (e.g., MobileNet (mobilenet, ), EfficientNet-Lite4 (efficientnet, ) DenseNet121 (densenet, ), and InceptionV2 (inception, )), $30$ models used for object detection and semantic segmentation tasks (e.g., RetinaNet (retinanet, ), Mask/Faster R-CNN R-50-FPN (mask-rcnn, ; faster-rcnn, ), and SSD (ssd, )), and $7$ transformer models (attention, ) utilized for text comprehension and generation tasks, such as question answering (Q&A), (BERT-SQuAD (BERT, )), sentiment analysis (RoBERTa (liu2020roberta, )), text summarization (T5 (T5Paper, )) and generic-purpose text generation (GPT-2 (gpt2, )), with $4$ models focusing on Q&A, while one model is utilized for each of the other text comprehension and generation tasks. We utilized models of a variety of opsets based on hub availability, ranging from $7$ (the lowest compatible with the latest ONNX Runtime version by the time of this publication - v1.21.11.21.1) up to $12$ (the highest provided from ONNX Model Hub via their API). Our methodology, however, is applicable to any ONNX model type and is not constrained by opset versions or tied to a specific model repository.
@@ 5.3. Computational Resources
We ran our models on mid-range systems with Intel i5 and i7 CPUs, reflecting the typical hardware used by developers during model development and fine-tuning. Our method is generalizable to larger models, limited only by ONNX model size constraints, and can be extended to high-end systems with a trivial extension for hardware acceleration (using the ONNX Runtime API (onnxruntime, ), or an AI compiler such as Apache TVM (tvm, )).
@@ 5.4. Datasets
We selected datasets appropriate to each model task. For classification and object detection, we used the ILSVRC2017 (ILSVRC17, ) validation dataset, containing $5500$ images, commonly used for benchmarking image recognition models (e.g., ResNet (resnet, ) and BERT (BERT, )). For text comprehension and generation, we used SQuAD-v2 (rajpurkar2018knowdontknowunanswerable, ), which includes $11873$ questions and contexts, to evaluate Q&A tasks and general-purpose transformer behavior. Additionally, for summarization and sentiment analysis, we used the IMDB dataset (moviedataset, ; imdbsentiment, ), containing $25000$ entries.
@@
  1. (1)
    For image recognition models performing classification tasks, we utilized the Kendall Rank Correlation Coefficient (kendall, ) (KT), measuring differences for top-K prediction labels of each dataset input image, for $K=1,5$, and $10$. If the KT value was found to be <1<1 for each input, we counted this input as different.
  2. (2)
    For image recognition models performing object detection and semantic segmentation, we calculated precision, recall, and F1 score for each dataset input, and extracted their average, combined with different thresholds of Intersection-over-Union (IoU) ($IoU=0.5,0.75$, and $0.9$), as well as IoU, mean Average Precision (mAP), along with average precision and recall. We considered the original model output as the base, and the optimizer model the one under test for our metric calculations, and the higher the values observed, the better the optimized model aligns with the original.
  3. (3)
    For text generation models performing question answering and summarization, we utilized BLEU score (BLEU, ) across output texts for each input, monitoring instances where BLEU_Score<1BLEU\_Score<1 for each input.
  4. (4)
    For sentiment analysis models, we compared each output value consisting of 0,1{0,1} across the original and the optimized versions, registering any differences observed.
@@ 6. Results
Following the execution of the full set of experiments. A complete list of these issues, along with the associated models, is shown in Table 2. The developers have acknowledged the issues we reported, and we have provided them with full reproducibility details for each reported bug. Of the $15$ bugs identified, $14$ were entirely new (not previously reported on the ONNX Optimizer issue tracker), while one was a previously reported issue that lacked sufficient detail, including model information and reproducibility steps. We provided comprehensive information and full reproducibility details for this issue.
Within the ONNX optimizer, the bugs identified were primarily associated with $9$ of the $47$ passes. However, not all passes are equally important. By default, the optimizer prioritizes passes related to fuse and eliminate operations. As a result, $8$ of the $11$ reported GitHub issues are considered more critical since they belong to the optimizer’s core (default) strategy. Among these $8$ issues, $6$ are linked to specific passes (e.g., fuse_bn_into_conv, which fuses batch normalization nodes with convolutions), while the remaining 2 issues were more general and related to the optimizer’s overall usage. We analyze and present the impact of these faults both on the optimization process and on the correctness of the optimized models in the following Sections.

### v5

@@ 3.5. Fault Localization
Once a method is selected, the Metrics Comparison Module calculates the metrics based on model type and generates a report with all results, in JSON format. If a difference is detected across the original and the optimized model, then DiTOX retrofits the reported data to the Optimizer Module. Then, the Optimizer Module generates model optimizations per-pass, reporting how each pass behaves. DiTOX applies a fault localization methodology for three cases: when the optimizer crashes, when it generates an optimized model in an invalid form, and when the optimization succeeds but produces results inconsistent with the original model. In addition, DiTOX registers any warnings that can indicate the presence of a potential performance fault (e.g., redundant initializers in the model). Upon detecting an issue or a problematic behavior, DiTOX attempts to identify the responsible pass by generating per-pass optimizations, executing inference on the resulting models, and comparing their outputs with those of the original model. This process is applied iteratively for each of the passes supported by the optimizer, until the pass responsible for the detected issue is identified.
@@ 4. Related Work
Table 1. Comparison of features between DiTOX and related work. Each technique is compared across the following parameters: testing of ONNX (ONNX?), ability for fault localization (Flt. Loc.), bug types, graph optimization passes (Gr. Opt. Passes), detection of accuracy changes (Acc.), crashes (Cr.), generation of malformed models due to miscompilation (Malf.), and issues affecting model execution times(T.) bugs, as well as the compilers explored on each contribution.
@@
DiTOX Per-pass fault localization, differential testing Logic, API versioning, graph input/parameters/ transformations Fuse/ eliminate/ rewrite ONNX Optimizer (40)
NNSmith (Liu et al., 2023) Grammar- based fuzzing †,‡ $\times$ Conversions, semantics, transformations, graph rewrites Conv2D inputs rewrite $\times$ $\times$ TVM (Chen et al., 2018), PyTorch JIT (Paszke et al., 2019), TensorRT (NVIDIA, 2019), PT2 (Ansel et al., 2024), ONNX Runtime (41), TF XLA (Sabne, 2020), TFLite (Abadi et al., 2016)
MLIRSmith (Wang et al., 2023a) Grammar- based fuzzing $\times$ $\times$ Conversions (dialects), generic transformations, logic (dialects) $\times$ $\times$ $\times$ $\times$ MLIR (Lattner et al., 2020b), LLVM (Lattner and Adve, 2004)
MLIRod (Suo et al., 2024) Operator dependency/ fuzzing $\times$ $\times$ Conversions (dialects), transformations $\times$ $\times$ $\times$ $\times$ MLIR (Lattner et al., 2020b)
@@
SYNTH- FUZZ (Limpanukorn et al., 2025) Grammar- based fuzzing $\times$ $\times$ Generic, input-based $\times$ $\times$ $\times$ $\times$ MLIR (Lattner et al., 2020b)
HirGen (Ma et al., 2023) Graph IR fuzzing $\times$ $\times$ Numeric computations, exceptions, types, memory $\times$ $\times$ $\times$ $\times$ TVM (Chen et al., 2018)
TVMFuzz (64) Low-level IR fuzzing $\times$ $\times$ Numeric computations, graph rewrites Constant folding $\times$ $\times$ TVM (Chen et al., 2018)
Tzer (Liu et al., 2022) Coverage- guided fuzzing (Low-level IR) $\times$ $\times$ API inconsistency/misuse, memory, types $\times$ $\times$ TVM (Chen et al., 2018)
GenCoG (Wang et al., 2023b) Constraint solving, High-level IR Fuzzing $\times$ $\times$ Memory, tensor shape /data, operator attributes, logic Operator fusion $\times$ $\times$ TVM (Chen et al., 2018)
DeepDiffer (Lin et al., 2023) Low-level IR/ coverage- guided fuzzing, differential testing $\times$ $\times$ Type, memory, API misuse, numeric computations, exception handling $\times$ $\times$ $\times$ TVM (Chen et al., 2018)
WhiteFox (Yang et al., 2024) LM analysis, fuzzing $\times$ $\times$ Memory, API misuse, concurrency, logic, security vulnerabilities $\times$ $\times$ TF XLA (Sabne, 2020), PyTorch Inductor (Paszke et al., 2019), LLVM (Lattner and Adve, 2004)
TorchProbe (Su et al., 2023) Feature- based fuzzing $\times$ $\times$ Memory, API misuse, logic $\times$ $\times$ $\times$ PyTorch (Paszke et al., 2019), Triton (Tillet et al., 2019)
MT- DLComp (Xiao et al., 2022) Metamorphic testing $\times$ $\times$ Memory, dead code, layout optimization, numeric accuracy Conv2D layout change $\times$ $\times$ TVM (Chen et al., 2018), Glow (Rotem et al., 2019), NNFusion (Ma et al., 2020), TF XLA (Sabne, 2020)
FetaFix (Louloudakis et al., 2025) Fault localiza- tion, differential testing $\times$ Preprocessing, input dimensions, parameters, hyperparameters, tensor shape, graph structure $\times$ $\times$ $\times$ $\times$ Converters across PyTorch (Paszke et al., 2019), Keras (Chollet, 2015), TF (Abadi et al., 2016), TFLite (Abadi et al., 2016)
In terms of the Technique column shown in Table 1, most existing works primarily rely on compiler fuzzing, often enhanced with techniques such as differential testing (e.g., DeepDiffer (Lin et al., 2023)) or LLM analysis (e.g., WhiteFox (Yang et al., 2024)). Consequently, these systems focus on triggering compiler bugs rather than pinpointing their origins and root causes, typically relying on generic compiler error messages to perform manual fault localization. Among these works, only FetaFix (Louloudakis et al., 2025) provides a mechanism for fault localization, but it targets accuracy-related faults in model conversion between DL frameworks and does not support the model optimization process. In contrast, DiTOX systematically localizes faults by applying an iterative fault localization methodology for each pass in the ONNX Optimizer.
In terms of bug types detected in Table 1, existing tools primarily focus on detecting crash bugs. For label accuracy-related bugs, they typically rely on random input generation (e.g., NNSmith (Liu et al., 2023)) to expose inconsistencies in model inference. In contrast, for label inconsistency bugs that DiTOX detects, we assess the prevalence and impact of these bugs on popular datasets(e.g., ILSVRC (Russakovsky et al., 2015), IMDB (Tripathi et al., 2020)) and model architectures. Additionally, only a small number of tools consider miscompilations (for instance, models that contain problematic input nodes or misconfigured parameters), with just two prior works capable of detecting faults of this kind (64; C. Yang, Y. Deng, R. Lu, J. Yao, J. Liu, R. Jabbarvand, and L. Zhang (2024)). The detection of bugs affecting execution times of the compiled models has received limited attention in the literature, with only two TVM-focused systems (Liu et al., 2022; Lin et al., 2023) using empirical measurements to evaluate performance-related faults. DiTOX currently identifies issues that could potentially lead to degraded model execution times by flagging redundant operator warnings (e.g., unused initializer entries).
Regarding support for ONNX in Table 1, DiTOX is the only tool capable of localizing faults in the ONNX Optimizer. Other methodologies either treat ONNX merely as an intermediate representation or focus on bug detection in ONNX Runtime (Yang et al., 2024), a cross-platform inference engine, unlike model optimization targeted by the ONNX Optimizer.
Existing work on bug detection within graph optimization passes (shown as Gr. Opt. Passes column in Table 1) is limited, with prior approaches focusing on common operations such as constant folding, operator fusion, and convolution optimizations. In contrast, DiTOX addresses the full spectrum of features in the ONNX Optimizer, covering a set of 47 passes that perform a wide variety of graph optimizations such as Fuse, Eliminate, and Rewrite operations in addition to the more common operations.
@@ 5.1. The Experiment Process
We evaluated all ONNX Optimizer passes (Section 2.3). For each model, we first applied the default setting, which runs all fusion and elimination passes, and compared the outputs of the original and optimized models as described in Section 3.4. When discrepancies arose, we re-optimized using individual passes, producing 47 single-pass models to identify the pass causing the issue. DiTOX also supports new or custom passes, and our methodology generalizes to other optimization frameworks (e.g., Apache TVM) that allow both bundled and individual pass execution.
@@ 5.2. DNN Models
For our experiment set, we sourced $130$ model instances from the ONNX Model Hub (39).333. The complete set of models used in our experiments, along with the full set of our results, is available in the results folder of the DiTOX repository at https://github.com/luludak/DiTOX/blob/master/results/README.md. In particular, we selected $93$ models utilized for classification tasks (e.g., MobileNetV2 (Howard et al., 2017), EfficientNet-Lite4 (10) DenseNet121 (Huang et al., 2018), as well as InceptionV2 (Szegedy et al., 2014b)), $30$ models used for object detection and semantic segmentation tasks (e.g., RetinaNet (Lin et al., 2020), Mask/Faster R-CNN R-50-FPN (He et al., 2020; Ren et al., 2015), and SSD (Liu et al., 2016)), and $7$ transformer models (Vaswani et al., 2017) utilized for text comprehension and generation tasks, such as question answering (Q&A), (BERT-SQuAD (Devlin et al., 2019)), sentiment analysis (RoBERTa (Liu et al., 2020)), text summarization (T5
 (Raffel et al., 2020)) and generic-purpose text generation (GPT-2 (21)), with $4$ models focusing on Q&A, while one model is utilized for each of the other text comprehension and generation tasks. We utilized models of a variety of opsets based on hub availability, ranging from $7$ (the lowest compatible with the latest ONNX Runtime version by the time of this publication - v1.21.11.21.1) up to $12$ (the highest provided from ONNX Model Hub via their API). Our methodology, however, is applicable to any ONNX model type and is not constrained by opset versions or tied to a specific model repository.
@@ 5.3. Computational Resources
We ran our models on mid-range systems with Intel i5 and i7 CPUs, reflecting the typical hardware used by developers during model development and fine-tuning. Our method is generalizable to larger models, limited only by ONNX model size constraints, and can be extended to high-end systems with a trivial extension for hardware acceleration (using the ONNX Runtime API (41), or an AI compiler such as Apache TVM (Chen et al., 2018)).
@@ 5.4. Datasets
We selected datasets appropriate to each model task. For classification and object detection, we used the ILSVRC2017 (Russakovsky et al., 2015) validation dataset, containing $5500$ images, commonly used for benchmarking image recognition models (e.g., ResNet (He et al., 2015) and BERT (Devlin et al., 2019)). For text comprehension and generation, we used SQuAD-v2 (Rajpurkar et al., 2018), which includes $11873$ questions and contexts, to evaluate Q&A tasks and general-purpose transformer behavior. Additionally, for summarization and sentiment analysis, we used the IMDB dataset (Maas et al., 2011; Tripathi et al., 2020), containing $25000$ entries.
@@
1. For image recognition models performing classification tasks, we utilized Kendall Rank Correlation Coefficient (Kendall, 1938) (KT), measuring differences for top-K prediction labels of each dataset input image, for $K=1,5$, and $10$. If the KT value was found to be <1<1, we treated this input as different.
2. For image recognition models performing object detection and semantic segmentation, we calculated precision, recall, and F1 score for each dataset input, and extracted their average, combined with different thresholds of Intersection-over-Union (IoU) ($IoU=0.5,0.75$, and $0.9$), as well as IoU, mean Average Precision (mAP), along with average precision and recall. We considered the original model output as the base, and the optimizer model the one under test for our metric calculations, and the higher the values observed, the better the optimized model aligns with the original.
3. For text generation models performing question answering and summarization, we utilized BLEU score (Papineni et al., 2002) across output texts for each input, monitoring instances where BLEU_Score<1BLEU\_Score<1 for each input.
4. For sentiment analysis models, we compared each output value consisting of 0,1{0,1} across the original and the optimized versions, registering any differences observed.
5. To distinguish actual faults from known optimizer limitations, we relied on ONNX Optimizer documentation and accounted for known limitations (e.g., split_init and split_predict are known to not fully comply with the ONNX standard format). Aside from issues from these two passes, all other detected behaviors were unexpected and thus classified as faults.
@@ 6. Results
Following the execution of the full set of experiments11footnotemark: 1, DiTOX successfully identified $15$ bugs, reported across $11$ separate GitHub issues (11) [^4] An overview of all detected issues using DiTOX can be found at https://github.com/luludak/DiTOX/blob/master/results/detected_issues.txt.. A complete list of these issues, along with the associated models, is shown in Table 2. The developers have acknowledged the issues we reported, and we have provided them with full reproducibility details for each reported bug. Of the $15$ bugs identified, $14$ were entirely new (not previously reported on the ONNX Optimizer issue tracker), while one was a previously reported issue that lacked sufficient detail, including model information and reproducibility steps. We provided comprehensive information and full reproducibility details for this issue.
Within the ONNX optimizer, the bugs identified were primarily associated with $9$ of the $47$ passes. However, not all passes are equally important. By default, the optimizer prioritizes passes related to fuse and eliminate operations. As a result, $8$ of the $11$ reported GitHub issues are considered more critical since they belong to the core (default) strategy of the optimizer. Among these $8$ issues, $6$ are linked to specific passes (e.g., fuse_bn_into_conv, which fuses batch normalization nodes with convolutions), while the remaining 2 issues were more general and related to the overall usage of the optimizer. We analyze and present the impact of these faults both on the optimization process and on the correctness of the optimized models in the following Sections.